Iza can earn 7800 for doing hours of office work and 2hrs of field work. However, she gets 8,200 if she works two hours in the office and 3 hours in the field. What is the rate per hour of Iza?


1
Expert's answer
2021-12-13T17:43:16-0500

Let us assume office work as ‘x’ and field work as ‘y’.

Given that, Iza zan earn 7800 for doing 3 hrs office work

So, 3x+2y = 7800 …(1)

Also, She gets 8200 for doing 2hrs office work and 3 hrs field work

2x+3y=8200 …(2)

Add (1) and (2), we get

5x+5y=16000

x+y=3200 …(3)

Subtracting (2) from (1), we get

x-y=-400 …(4)

Add (3) and (4), we get

2x=2800

x=1400

From (3),

1400+y=3200

y = 1800

Therefore, she gets 1400 for 4 hr office work and 1800 for 1 hr field work.